Grade 11Physical Chemistry

How many chlorine atoms can be ionized in the process?
Cl------> Cl+ + e- by the energy liberated from the following process:
Cl + e -----> Cl- for 6 x1023atoms
given that electron affinity of chlorine is 3.61 eV and ionization energy of chlorine is 17.422 eV.

Meera
10 Years agoGrade 11

2 Answers

Sakshi
10 Years ago
hey student..
for this question, you need to take care of the units of the electron affinity values.
try to decode the defination given in the question and go one by one to both the definations and you ll be getting the required answer
Divya
10 Years ago
The answer given is 1.24 * 10^23 .. from working it out, my answer was 124 * 10^23.. :
energy = -13.6 z^2/n^2 where z in atomic number and n is the shell number. so z for chlorine is 17 and so z^2 is 289. Writing down the electronic configuration, 1s2 2s2 2p6 3s2 3p2 so n is 3. and n^2 is 9. so energy produced is 435.2 eV. so x(17.422+3.61) = 435.2 * 6 * 10^23. so x = 124* 10^23.. I am not able to spot my mistake... or maybe the answer given is wrong..